# NaviAgent: Bilevel Planning on Tool Dependency Graphs for Function Calling

### v1

@@ Abstract
While large language models (LLMs) have demonstrated remarkable potential in open-domain reasoning, their reliance on static knowledge and fragile tool invocation severely hinder the orchestration of complex, heterogeneous toolchains, especially when scaled to thousands of tools. Existing methods typically use rigid single-path execution, resulting in poor error recovery and exponentially growing search spaces. We introduce NaviAgent, a graph-navigated bilevel planning architecture for robust function calling, comprising a Multi-Path Decider and Graph-Encoded Navigator. As an LLM-powered agent, the Multi-Path Decider defines a four-dimensional decision space and continuously perceives environmental states, dynamically selecting the optimal action to fully cover all tool invocation scenarios. The Graph-Encoded Navigator constructs a Tool Dependency Heterogeneous Graph (TDHG), where node embeddings explicitly fuse API schema structure with historical invocation behavior. It also integrates a novel heuristic search strategy that guides the Decider toward efficient and highly successful toolchains, even for unseen tool combinations. Experiments show that NaviAgent consistently achieves the highest task success rate (TSR) across all foundation models and task complexities, outperforming the average baselines (ReAct, ToolLLM, $\alpha$-UMI) by 13.5%, 16.4%, and 19.0% on Qwen2.5-14B, Qwen2.5-32B, and Deepseek-V3, respectively. Its execution steps are typically within one step of the most efficient baseline, ensuring a strong balance between quality and efficiency. Notably, a fine-tuned Qwen2.5-14B model achieves a TSR of 49.5%, surpassing the much larger 32B model (44.9%) under our architecture. Incorporating the Graph-Encoded Navigator further boosts TSR by an average of 2.4 points, with gains up over 9 points on complex tasks for larger models (Deepseek-V3 and GPT-4o), highlighting its essential role in toolchain orchestration.
@@ 1 Introduction
Integration of large language models (LLMs) with external tools has evolved from single-tool interactions parisi2022talm ; schick2023toolformer ; chen2024towards to function calling agents shen2023hugginggpt ; yang2023gpt4tools capable of executing complex, multi-stage procedures qu2025tool . However, critical real-world challenges such as API instability, rigid workflow specifications remain unresolved. Despite incremental improvements from Chameleon’s modular heterogeneous tool integration lu2023chameleon , ToolVerifier’s self-verification paradigm mekala2024toolverifier , and Confucius’s iterative self-guidance mechanism gao2024confucius , all such methods ultimately inherit the weakness of static workflow architectures maldonado2024multi ; krishnan2025ai - an inability to dynamically adapt to runtime contingencies. When critical API services are unavailable, they lack mechanisms for dynamic fallback tran2025multi ; jiang2025appagentx or alternative path discovery, severely limiting real-world deployment reliability, especially in complex multi-tool chaining call scenarios barua2024exploring ; shi2024chain .
Compounding these challenges, scalability issues escalate dramatically as tool libraries grow. Heuristic approaches like ToolLLM qin2023toolllm and ToolChain* zhuang2023toolchain reduce search complexity through functional hierarchy trees and A* cost estimation, yet remain fundamentally constrained by static tool relationship assumptions. While ToolNet liu2024toolnet attempts to address this via dynamic dependency graphs constructed from historical invocation logs, its practical utility is limited by sparse multi-hop interactions, only 12% of APIs in production logs exhibit more than 3 chained calls, leaving long-tail tools disconnected. This sparsity creates a cold-start bottleneck zhang2025cold ; zhang2024llmtreerec , raising a core question: Can we automatically infer potential tool collaboration patterns by combining inherent structural features (e.g., parameter types, I/O constraints) rather than relying solely on historical data?
Building upon the graph-enhanced LLM paradigm wu2024can , we present NaviAgent, a novel graph-guided framework for robust function calling. NaviAgent features a bilevel planning architecture that integrates a Multi-Path Decider and a Graph-Encoded Navigator, enabling dynamic toolchain orchestration and error recovery. The Multi-Path Decider defines a four-dimensional decision space and continuously perceives environmental states to flexibly select optimal actions for both routine and exceptional scenarios. The Graph-Encoded Navigator, with a novel heuristic search strategy, constructs and navigates a Tool Dependency Heterogeneous Graph (TDHG) that captures both structural and behavioral relationships among tools. This closed-loop mechanism integrates graph-based reasoning with real-time execution feedback, allowing the agent to discover alternative paths and adapt to runtime contingencies. As a result, NaviAgent robustly handles both simple and complex tool invocation scenarios, supporting scalable and reliable deployment in large, heterogeneous tool environments. Our experiments show that NaviAgent substantially outperforms the strongest baseline ($\alpha$-UMI) on Deepseek-V3, achieving improvements of 15.4% in task completion rate and 14.8% in task success rate (reaching 98.3% and 55.5%, respectively). The integration of the Graph-Encoded Navigator consistently provides a average 2.4 point boost over non-graph baselines across all tasks, with gains of up to 8 points on complex tasks for larger models, underscoring its critical role in complex toolchain coordination. Furthermore, our framework enables a fine-tuned lightweight Qwen2.5-14B model to even surpass the much larger Qwen2.5-32B model in task success rate (49.5% vs. 44.9%), demonstrating NaviAgent’s efficiency and scalability.
Our main contributions are as follows: i) NaviAgent Architecture: A scalable bilevel planning framework that orchestrates toolchains at thousand-scale while achieving superior performance to SOTA baselines, maintaining an optimal quality-efficiency equilibrium. ii) Multi-Path Decider: A four-dimensional decision space covering the full spectrum of tool invocation scenarios. iii) Dynamic Graph-Encoded Navigator: Joint learning of node features and adaptive edge weights enables continuous optimization and discovery of both existing and novel tool combinations.
@@ 2 Related Work
Early research focused on enhancing LLMs’ single-tool invocation capabilities. TALM parisi2022talm established foundational paradigms through predefined template chains, while EasyTool yuan2024easytool introduced structured tool descriptions to reduce semantic parsing overhead. For long-context scenarios, tool documentation compression techniques preserved critical semantics via summarization, enabling low-resource tool usage xu2024concise . Toolformer schick2023toolformer innovatively embedded tool invocation APIs in pre-training, allowing self-supervised learning of usage patterns from unlabeled data. In multimodal settings, GPT4Tools yang2023gpt4tools improved visual tool generalization (e.g. object detection) by aligning vision-language instructions with tool descriptions.
@@
As tool libraries expanded, HuggingGPT shen2023hugginggpt proposed a four-stage pipeline (plan, select, execute, respond) for standardized multi-tool workflows, while Chameleon lu2023chameleon integrated heterogeneous tools (13+ types) via modular composition. Similarly, $\alpha$-UMI shen2024small decomposes the tool-use process into planning, invocation, and summarization, but uniquely assigns each stage to a dedicated lightweight LLM, enabling modular updates and improved performance, especially for smaller models. For small toolkits, TRICE qiao2023making optimized single tool policies via execution feedback, and ToolFactory ni2025toolfactory automated tool adaptation through domain-guided code synthesis. However, these approaches struggled with dynamic collaboration. For large-scale toolkits, Confucius gao2024confucius addressed combinatorial explosion via hierarchical tool classification, and ToolVerifier mekala2024toolverifier improved selection robustness through self-verification mechanisms.
@@
Static frameworks faltered under open-domain task complexity, prompting dynamic decision mechanisms. ReAct yao2023react pioneered the decoupling of reasoning from tool calls through chain-of-thought planning. Building on this, Reflexion shinn2023reflexion enhanced error recovery by introducing iterative self-reflection, significantly improving fault tolerance in complex tasks. For long-horizon tasks, path search techniques became pivotal: Tree-of-Thoughts (ToT) yao2023tree formalized tool invocation as searchable reasoning trees with dynamic branching, while ToolLLM qin2023toolllm optimized search efficiency through functional hierarchy-guided DFS. ToolChain zhuang2023toolchain further advanced this by employing heuristic cost estimation to prioritize high-success-rate branches. Yet, these methods assumed static tool relationships, failing to adapt to API drift or cross-domain tasks. ControlLLM liu2024controlllm built static dependency graphs for task decomposition, whereas ToolNet liu2024toolnet dynamically updated tool relations from historical calls, both limited by sparse multi-hop interaction data. This gap motivates our dynamic Graph-Encoded Navigator that jointly models structural dependencies and behavioral adaptations to capture evolving tool relations, aligning with findings that graph learning enhance LLM planning wu2024can ; besta2024graph .
@@ 3 Methodology
Figure 1: The overview of NaviAgent.
Figure 1 presents the NaviAgent architecture, which integrates a Multi-Path Decider and a dynamically updating Graph-Encoded Navigator in a bilevel planning framework. The Decider, shown in the blue box, selects among actions such as direct response, intent clarification, tool retrieval, or tool call. The Navigator, visualized as five dark brown components, is responsible for planning toolchains over an evolving TDHG with learned edge weights, as depicted in the graph structure and its temporal evolution. As the Decider interacts with the environment, the graph structure and edge weights are continuously updated based on execution feedback, enabling adaptive and robust planning. The bottom of the figure demonstrates NaviAgent’s ability to learn more optimal routing strategies over time, as well as to flexibly adapt its execution path in response to API unavailability. Further details on the graph evolution mechanism are provided in Section 3.2.4.
3.1 Multi-Path Decider
3.1.1 LLM-Driven Autonomous Agent Architecture
[…]
@@ 3.1.1 LLM-Driven Autonomous Agent Architecture
where a sliding window maintains the most recent three [^1] Our experiments demonstrate that utilizing the most recent three observation-action pairs achieves the best balance between accuracy and efficiency. observation-action pairs, capturing the Decider’s recent decision trajectory. The pruned tool dependency subgraph $\mathcal{G}_{t-1}^{\prime}=(V,E,W)$ is computed based on the Decider’s state at the previous time step $t-1$, where $V$ is the node set, $E$ is the edge set, and $W$ denotes the edge weights indicating dependency strengths. The overall decision function is then formulated as
@@ 3.1.2 Model Training
For supervised fine-tuning, we adopt the standard language modeling objective, computing the loss exclusively over the response or action generation segments. During training, the LLM-based decider receives as input the most recent historical state-action pairs $\mathcal{H}_{t}$, the current observation $\mathcal{O}_{t}$, and the pruned tool dependency subgraph $\mathcal{G}_{\mathrm{sub}}$. The model is trained to maximize the likelihood of the ground-truth action $a_{t}^{*}$ at step $t$, which is derived from high-quality, curated datasets (see Appendix E.2 for details):
@@
where $N$ is the number of training samples and $p_{\theta}$ denotes the Decider’s predicted probability over the action space.
3.2 Graph-Encoded Navigator
@@ 3.2.1 Graph Construction and Representation
We construct a TDHG, a directed weighted graph $\mathcal{G}=(V,E,W)$ with API and parameter nodes. Edges include structural chains, defined by API schemas (e.g., parameter-to-API and API-to-parameter connections), as well as behavioral chains, derived from historical usage data (e.g., API-to-API and parameter-to-parameter dependencies). Each edge is assigned a statistical weight $\tilde{w}_{ij}$ reflecting empirical invocation patterns.
@@ 3.2.2 Graph Training Objective
The graph model is trained with a hybrid loss that combines cross-entropy $\mathcal{L}_{CE}$ and adaptive margin $\mathcal{L}_{\mathrm{Margin}}$ objectives (see Appendix B.2 for details), both leveraging edge weights $\tilde{w}_{uv}$ to capture graded dependencies. Specifically, statistical edge weights serve as soft labels in the cross-entropy loss and dynamically adjust the margin in the margin loss, enabling the model to focus on more critical dependencies. The final objective is a weighted sum of these two terms:
@@
where $\mu_{t}=\mu_{0}\cdot\gamma^{t}$ is a decaying weight over training epochs $t$, with initial value $\mu_{0}$ and decay rate $\gamma\in(0,1)$. This curriculum-inspired strategy first emphasizes classification accuracy and then prioritizes representation discrimination, leading to both robust prediction and well-structured embeddings.
@@ 3.2.3 Graph Search
At inference time, the predicted link probabilities $p_{uv}$ are assigned as edge weights $w_{uv}$ in the TDHG, providing the foundation for graph search and toolchain planning. We implement two classes of search strategies.
Alpha-Beta Pruning.
This algorithm knuth1975analysis is adapted for backward search over the TDHG, using dynamic thresholding and path scoring to prune suboptimal toolchains. The core scoring function balances direct edge weights, target connectivity, and path depth:
where $w_{u\rightarrow v_{t}^{\mathrm{api}}}$ and $w_{u\rightarrow v_{t}^{\mathrm{param}}}$ denote the edge weights from $u$ to the target API node $v_{t}^{\mathrm{api}}$ and target parameter node $v_{t}^{\mathrm{param}}$, respectively, included only if the corresponding indicator function $\mathbb{I}(\cdot)$ is active. And $\mathcal{D}(d)$ penalizes longer paths.
Heuristic Graph Search with Dynamic Pruning.
Our hybrid heuristic search algorithm combines simulated annealing kirkpatrick1983optimization and genetic algorithm strategies shapiro1999genetic , using a composite fitness function to evaluate candidate toolchains:
where $\mathcal{C}_{c}$ (node compactness) measures the closeness centrality of API nodes, $\rho_{p}$ (parameter density) is the ratio of parameter nodes within the subgraph to promote concise yet informative solutions, $\mathcal{D}_{c}=0.2e^{-d/10}+0.8e^{-n/8}$ (depth penalty) penalizes overly deep or complex dependency structures, with $d$ as the average depth and $n$ as the total node count, $\mathcal{W}_{n}$ (weight quantification) encourages solutions with higher cumulative edge weights, and $\mathcal{C}_{p}$ (path complexity) evaluates structural simplicity, favoring solutions with less intricate connectivity. Complete parameter settings, algorithmic details for both search strategies are provided in Appendix B.3 and Algorithm 1,  2).

### v2

# NaviAgent: Bilevel Planning on Tool Navigation Graph for Large-Scale Orchestration

@@ Abstract
Large language models (LLMs) have recently demonstrated the ability to act as function call agents by invoking external tools, enabling them to solve tasks beyond their static knowledge. However, existing agents typically call tools step by step at a time without a global view of task structure. As tools depend on each other, this leads to error accumulation and limited scalability, particularly when scaling to thousands of tools. To address these limitations, we propose NaviAgent, a novel bilevel architecture that decouples task planning from tool execution through graph-based modeling of the tool ecosystem. At the task-planning level, the LLM-based agent decides whether to respond directly, clarify user intent, invoke a toolchain, or execute tool outputs, ensuring broad coverage of interaction scenarios independent of inter-tool complexity. At the execution level, a continuously evolving Tool World Navigation Model (TWNM) encodes structural and behavioral relations among tools, guiding the agent to generate scalable and robust invocation sequences. By incorporating feedback from real tool interactions, NaviAgent supports closed‑loop optimization of planning and execution, moving beyond tool calling toward adaptive navigation of large‑scale tool ecosystems. Experiments show that NaviAgent achieves the best task success rates across models and tasks, and integrating TWMN further boosts performance by up to 17 points on complex tasks, underscoring its key role in toolchain orchestration.
@@ 1 Introduction
Large language models (LLMs) are increasingly deployed as function call agents, moving beyond single utilities toward complex multi‑stage workflows shen2023hugginggpt ; yang2023gpt4tools ; qu2025tool . However, real‑world environments contain thousands of heterogeneous tools that are continually updated, while tasks demand long sequences of coordinated invocations. Agents built around fixed tool descriptions or rigid workflows fail to adapt, making API drift, continual updates, and unseen tool compositions key challenges for function call agents.
Existing approaches attempt to mitigate brittleness but remain incomplete. Some embed tool knowledge directly into model parameters wang2410toolgen , which reduces context demands but requires costly retraining when APIs change. Others derive static graphs from invocation logs liu2024toolnet , yet sparse traces and missing parameter relations hinder generalization. Policy‑adaptation methods adjust individual tools with feedback chen2024learning , while clustering‑based planners enable substitutions liu2024tool . Taken together, existing methods can be broadly categorized into two camps: either structured but static, failing to evolve with the ecosystem, or adaptive but unstructured, lacking the representations needed to capture composability and complementarity.
Underlying these challenges is the complexity of the tool ecosystem: it spans thousands of heterogeneous tools, exhibits interdependencies such as parameter flows and functional complementarity, and evolves continually through addition, update, and deprecation. Such properties reveal why step‑by‑step invocation without global awareness cannot achieve reliable tool composition. The difficulty is compounded by the fact that API documentation is written for humans and often misaligns with how models interpret and use individual tools qu2024exploration , while flat catalogues provide little information on how tools compose, substitute, or adapt as the ecosystem changes. What is needed is a structured representation learned from execution traces that makes these dependencies explicit and continually adapts with feedback.
We propose NaviAgent, a bilevel planning framework that decouples high‑level task reasoning from low‑level execution. At the planning level, NaviAgent defines a four-dimensional decision space (direct response, intent clarification, toolchain retrieval, tool execution) covering core tool invocation scenarios, allowing the agent to operate without reasoning over complex inter-tool connections. At the execution level, it constructs the Tool World Navigation Model (TWNM), which encodes both structural and behavioral dependencies learned from execution traces. By coupling these graph‑based representations with navigation strategies, TWNM enables retrieval, substitution, and multi‑tool composition as the ecosystem evolves. Execution feedback continually updates both TWNM and the decision policy, forming a closed loop for robust adaptation to changing APIs.
Our main contributions are as follows: i) NaviAgent Architecture. The first bilevel agent framework that decouples high‑level task planning from low‑level tool execution, enabling scalable task composition across thousands of tools while preserving efficiency. ii) Tool World Navigation Model. A unified model that captures inter‑tool structures and behavioral dependencies from execution traces, and supports navigation and flexible search in large‑scale tool ecosystems. iii) Closed‑loop Evolution. A feedback mechanism where execution traces continuously refine TWNM and decision strategies, driving the co‑evolution of representation and decision‑making.
@@ 2 Related Work
Early research focused on enhancing LLMs’ single-tool invocation capabilities. TALM (parisi2022talm, ) established foundational paradigms through predefined template chains, while EasyTool (yuan2024easytool, ) introduced structured tool descriptions to reduce semantic parsing overhead. For long-context scenarios, tool documentation compression techniques preserved critical semantics via summarization, enabling low-resource tool usage (xu2024concise, ). Toolformer (schick2023toolformer, ) innovatively embedded tool invocation APIs in pre-training, allowing self-supervised learning of usage patterns from unlabeled data. In multimodal settings, GPT4Tools (yang2023gpt4tools, ) improved visual tool generalization (e.g. object detection) by aligning vision-language instructions with tool descriptions.
@@
As tool libraries expanded, HuggingGPT (shen2023hugginggpt, ) proposed a four-stage pipeline (plan, select, execute, respond) for standardized multi-tool workflows, while Chameleon (lu2023chameleon, ) integrated heterogeneous tools (13+ types) via modular composition. Similarly, $\alpha$-UMI (shen2024small, ) decomposes the tool-use process into planning, invocation, and summarization, but uniquely assigns each stage to a dedicated lightweight LLM, enabling modular updates and improved performance, especially for smaller models. For small toolkits, TRICE (qiao2023making, ) optimized single tool policies via execution feedback, and ToolFactory (ni2025toolfactory, ) automated tool adaptation through domain-guided code synthesis. However, these approaches struggled with dynamic collaboration. For large-scale toolkits, Confucius (gao2024confucius, ) addressed combinatorial explosion via hierarchical tool classification, and ToolVerifier (mekala2024toolverifier, ) improved selection robustness through self-verification mechanisms.
@@
Static frameworks faltered under open-domain task complexity, prompting dynamic decision mechanisms. ReAct (yao2023react, ) pioneered the decoupling of reasoning from tool calls through chain-of-thought planning. Building on this, Reflexion (shinn2023reflexion, ) enhanced error recovery by introducing iterative self-reflection, significantly improving fault tolerance in complex tasks. For long-horizon tasks, path search techniques became pivotal: Tree-of-Thoughts (ToT) (yao2023tree, ) formalized tool invocation as searchable reasoning trees with dynamic branching, while ToolLLM (qin2023toolllm, ) optimized search efficiency through functional hierarchy-guided DFS. ToolChain (zhuang2023toolchain, ) further advanced this by employing heuristic cost estimation to prioritize high-success-rate branches. Yet, these methods assumed static tool relationships, failing to adapt to API drift or cross-domain tasks. ControlLLM (liu2024controlllm, ) built static dependency graphs for task decomposition, whereas ToolNet (liu2024toolnet, ) dynamically updated tool relations from historical calls, both limited by sparse multi-hop interaction data. This gap motivates our TWNM that jointly models structural dependencies and behavioral adaptations to capture evolving tool relations, aligning with findings that graph learning enhance LLM planning (wu2024can, ; besta2024graph, ).
@@ 3 Methodology
Figure 1: Conventional function call agents vs. Our NaviAgent.
A key challenge in function call agents is that API calls usually have interdependent parameters and strict invocation orders. NaviAgent addresses this challenge by retrieving the entire toolchain before execution, rather than calling APIs step by step. This global reasoning allows the agent to plan an end-to-end path from the user query to the target API once and execute it directly. As illustrated in Figure 1, NaviAgent avoids repeated retrievals and automatically discards unreliable APIs, leading to more efficient and robust execution.
3.1 A Four-dimensional Decision Agent
3.1.1 Definition
The architecture achieves end-to-end decision-making through LLMs, formally modeled as a quintuple $(\mathcal{H},\mathcal{O},\mathcal{G},\mathcal{A},F)$ where $\mathcal{H}=\{(o_{t-i},a_{t-i})\}_{i=1}^{n}$ represents historical states (containing state sequence $\{o_{i}\}$ and action sequence $\{a_{i}\}$), $\mathcal{O}$ denotes the observation, $\mathcal{G}$ represents the tool dependency graph, $\mathcal{A}=\{\emph{Direct Response},\ \emph{Intent Clarification},\ \emph{ToolChain Retrieval},\ \emph{Tool Execution}\}$ defines the four dimensional decision space, where each action corresponds to directly answering the user, requesting clarification, retrieving candidate tool dependency subgraph via TWNM, or execute selected toolchains, respectively. $F:\mathcal{H}\times\mathcal{O}\times\mathcal{G}\to\mathcal{A}$ specifies the decision function. At each time step $t$, the agent constructs its decision context as follows. The historical context $\mathcal{H}_{t}$ is defined as
@@ 3.1.1 Definition
where a sliding window maintains the most recent three [^1] Our experiments demonstrate that utilizing the most recent three observation-action pairs achieves the best balance between accuracy and efficiency. observation-action pairs, capturing the agent’s recent decision trajectory. The pruned tool dependency subgraph $\mathcal{G}_{t-1}^{\prime}=(V,E,W)$ is computed from the agent’s state at the previous time step t1t-1, where $V$ is the node set, $E$ is the edge set, and $W$ denotes the edge weights indicating dependency strengths. The subgraph is serialized into a tree-structured textual format, ensuring a simplified yet sufficient representation for selected toolchains. The overall decision function is then formulated as
@@ 3.1.2 Model Training
For supervised fine-tuning, we adopt the standard language modeling objective, computing the loss exclusively over the response or action generation segments. During training, the LLM-based agent receives as input the most recent historical state-action pairs $\mathcal{H}_{t}$, the current observation $\mathcal{O}_{t}$, and the pruned tool dependency subgraph $\mathcal{G}_{\mathrm{sub}}$. The model is trained to maximize the likelihood of the ground-truth action $a_{t}^{*}$ at step $t$, which is derived from high-quality, curated datasets (see Appendix E.2 for details):
@@
where $N$ is the number of training samples and $p_{\theta}$ denotes the agent’s predicted probability over the action space.
3.2 Tool World Navigation Model
Figure 2: Tool dependency graph and its temporal evolution in TWNM. The left part shows the overall dependency relations, while the right part illustrates the pruning and evolution of executable subgraphs across time steps.
@@ 3.2.1 Graph Construction and Representation
We construct a directed weighted graph $\mathcal{G}=(V,E,W)$ with API and parameter nodes. Edges include structural chains, defined by API schemas (e.g., parameter-to-API and API-to-parameter connections), as well as behavioral chains, derived from historical usage data (e.g., API-to-API and parameter-to-parameter dependencies) , as illustrated in Figure 2 (left). Each edge is assigned a statistical weight $\tilde{w}_{ij}$ reflecting empirical invocation patterns.
@@ 3.2.2 Graph Training Objective
The graph model is trained with a hybrid loss that combines cross-entropy and adaptive margin objectives, both leveraging edge weights $\tilde{w}_{uv}$ to capture graded dependencies.
Cross‑entropy.
Adaptive margin.
It assigns larger separation to higher‑weight edges(i.e., $\tilde{w}_{uv}\rightarrow 1$), focusing learning on critical dependencies. For each positive edge $(u,v)^{+}\in\mathcal{E}^{+}$, $k$ negative edges $\{(u_{j},v)\}_{j=1}^{k}$ are sampled to construct positive and negative pairs for the margin loss.
$m_{uv}=m_{0}\left(1+\sigma(\tilde{w}_{uv})\right)$ (7)
where $m_{0}$ is a base margin, $\sigma(\cdot)$ denotes the sigmoid function, $\tilde{w}_{uv}$ is the statistical edge weight, $s(u,v)$ measures the embedding similarity, $(u,v)^{+}$ represents a positive edge, $(u_{j},v)^{-}$ denotes a negative sample, and $\cdot$$_{+}$ is the hinge function, and $\mathcal{E}^{+}$ is the set of positive edges.
The final training objective is a weighted sum of the two losses:
$\mu_{t}=\mu_{0}\cdot\gamma^{t},\ \gamma\in(0,1)$ (9)
@@
where $\mu_{t}$ is the weight for the cross-entropy loss at epoch $t$, $\mu_{0}$ is the initial weight, and $\gamma$ is a decay factor controlling the rate at which the contribution of the cross-entropy loss decreases over training. This curriculum strategy first emphasizes accuracy, then discrimination, yielding accurate predictions and structured embeddings.
@@ 3.2.3 Graph Search
At inference time, the predicted link probabilities $p_{uv}$ are used as edge weights $w_{uv}$ in the tool graph, forming the basis for weighted-graph search and toolchain planning. We adopt two representative search strategies adapted to this setting: an Alpha-Beta Pruning method that eliminates weak toolchains using dynamic thresholds, and a heuristic search that evaluates candidate toolchains with a composite fitness balancing connectivity, depth, and cumulative weights. Complete algorithms and parameter details are provided in Appendix B.2.